OpenStudy (anonymous):

What number would you have to add to both sides to complete the square? x2 – 14x = 7 A. 49 B. –196 C. 196 D. –49

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

take half of x coefficient square it

OpenStudy (anonymous):

49

ganeshie8 (ganeshie8):

\(x^2 – \color{red}{14}x = 7 \)

ganeshie8 (ganeshie8):

yes, 14/2 = 7 7^2 = 49

ganeshie8 (ganeshie8):

so you need to add 49 both sides to complete the square